### v2

# Noncooperative Equilibrium Selection via a Trading-based Auction

Abstract
Noncooperative multi-agent systems often face coordination challenges due to conflicting preferences among agents. In particular, agents acting in their own self-interest can settle on different equilibria, leading to suboptimal outcomes or even safety concerns. We propose an algorithm named trading auction for consensus (TACo), a decentralized approach that enables noncooperative agents to reach consensus without communicating directly or disclosing private valuations. TACo facilitates coordination through a structured trading-based auction, where agents iteratively select choices of interest and provably reach an agreement within an a priori bounded number of steps. A series of numerical experiments validate that the termination guarantees of TACo hold in practice, and show that TACo achieves a median performance that minimizes the total cost across all agents, while allocating resources significantly more fairly than baseline approaches.
{IEEEkeywords}
Decentralized control, Equilibrium selection, Noncooperative games, Multi-agent systems
1 Introduction
Multi-agent systems often require consensus among agents, yet such consensus can be challenging due to conflicts and disagreements. These conflicts stem from differences in the strategic priorities of the agents, making coordination difficult without communication or external mediation. A key example is the equilibrium selection problem [1], where agents must choose one outcome from multiple Nash equilibria—stable outcomes where no agent can unilaterally improve their payoff. Without a mechanism to enforce a consensus, these disagreements can cause suboptimal outcomes and even safety risks in critical applications [2, 3].
Figure 1: Illustration of a waypoint merging scenario in air traffic management, where multiple aircraft must coordinate their estimated time of arrival (ETAs) at a shared waypoint. Each equilibrium represents a possible sequence of arrival times that ensures safe time separation between aircraft, highlighting the need for consensus to avoid conflicts and inefficiencies.
Reaching a consensus in multi-agent systems is particularly difficult when agents are noncooperative, and this difficulty arises from several factors. First, the consensus process must be procedurally rational, meaning noncooperative agents should have an incentive to participate. Second, in practical cases, centralized coordination may be unavailable, necessitating decentralized approaches. Additionally, noncooperative agents often avoid sharing private information and lack access to 1-on-1 communication, further complicating coordination.
Existing consensus methods often rely on centralized coordination [4, 5, 6, 7, 8, 9], cooperation assumptions [10, 11, 1, 12], or direct communication [13, 14, 15], limiting their applicability for noncooperative scenarios. For example, voting—a common mechanism for achieving agreement—often leads to outcomes that some agents reject [16]. Therefore, we need a decentralized approach that achieves agreement over conflicting choices, respects individual preferences, and preserves privacy without requiring direct communication.
We propose an algorithm called the trading auction for consensus (TACo) that enables agents to reach an agreement in noncooperative situations. This algorithm is a structured trading-based auction where agents iteratively select choices based on self-interest. Over multiple rounds, agents attempt to persuade others to adopt their preferred choice by offering trades. When disagreements persist, they refine their negotiation strategies, making increasingly precise adjustments to their offers. The process continues until no agent sees a meaningful benefit in further negotiation, ensuring a stable outcome—a property that is provably satisfied within a finite number of steps.
TACo addresses the core challenges of noncooperative consensus by enabling agents to adjust their preferences based on individual benefit, while ensuring decentralization and privacy preservation. It preserves privacy by not disclosing valuations over choices or trading assets. The algorithm does not require a centralized coordinator and uses broadcast communication—a common method of sharing information in multi-agent systems—which avoids direct communication. This comprehensive design makes TACo applicable to a range of noncooperative multi-agent systems, including autonomous vehicles and coordinating drone operations in cities.
The contributions of this work are threefold: (i) We introduce the trading auction for consensus (TACo) algorithm, which enables consensus in noncooperative settings without requiring a centralized coordinator or direct communication, and while preserving privacy. (ii) We prove that TACo enforces an agreement in finite time, and derive an explicit upper bound on the number of steps required for all agents to reach this mutual agreement. (iii) We validate our theoretical findings through numercial experiments and demonstrate that TACo outperforms existing negotiation methods regarding social optimality and fairness in a practical equilibrium selection scenario.
2 Related Work
Existing multi-agent consensus methods employ leader-follower frameworks [17, 18], finite-time consensus algorithms [19, 20], robust control techniques [21], auction mechanisms [22], and game-theoretic approaches for noncooperative agents [23]. These have been applied in UAV coordination [24, 25, 26], traffic flow optimization [23], and wireless sensor networks [27]. However, most rely on centralized leaders or high-bandwidth communication, limiting their applicability in fully decentralized, noncooperative settings [28].
A similar trend appears in addressing the equilibrium selection problem, where agents must coordinate among multiple conflicting outcomes. Many approaches rely on a centralized coordinator, requiring agents to share private information or follow the coordinator’s directives [4, 5, 6, 13, 14, 7, 15, 8, 9]. Others assume cooperativeness or shared beliefs, which bypass individual preferences [10, 11, 1, 12]. While effective in achieving a desirable equilibrium, these approaches are less suited for fully noncooperative systems due to their reliance on direct communication and lack of procedural rationality, where agents have incentives to participate based on self-interest.
Several works tackle equilibrium selection in noncooperative systems [12, 1, 29, 10, 11, 30, 31]. These include enforcing unique equilibria by modifying utility functions [29] or employing structured processes like the Linear Tracing Procedure, which relies on shared initial beliefs [10, 11]. While ensuring rationality during equilibrium seeking, these methods assume a coordinator or shared preferences during equilibrium selection.
Negotiation algorithms provide another approach to decentralized consensus among noncooperative agents [16]. Common methods include voting, bargaining, and auction-based mechanisms. Voting is widely used but often fails to ensure rationality, meaning outcomes may not be acceptable to all agents [16]. Bargaining relies on direct communication and is often problem-specific [32, 33, 34], limiting its applicability in general decentralized negotiation environments. In contrast, auctions align agents’ incentives with self-interest while preserving privacy (i.e., agents need not disclose sensitive private information) [35, 32, 36, 16, 37]. However, auctions primarily address resource allocation rather than single-choice problems like equilibrium selection, highlighting the need for research that bridges this gap by adapting auction-based methods for decentralized, noncooperative equilibrium selection.
3 Coordination Under Conflicting Preferences
A prominent instance of conflicting choices in multi-agent systems is the equilibrium selection problem. Before explaining the problem, we first introduce the notation and framework used to model the interactions between agents.
3.1 Equilibrium selection problem
We model the interaction between agents as a noncooperative game with $n\in\mathbb{Z}_{\geq 0}$ agents, where each agent $i\in[n]\equiv\{1,2,\ldots,n\}$ selects an action $x_{i}$ from an action set $\mathbb{X}_{i}$. The collective actions of all agents are represented by the joint action profile $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{n}\}$. Each agent $i$ wishes to minimize a cost function $c_{i}(x_{i},\mathbf{x}_{\neg i})$ which depends on its own action $x_{i}$ and those of other agents, where $\mathbf{x}_{\neg i}=\mathbf{x}\setminus x_{i}$. The collection of costs of all agents is denoted as $\mathbf{c}=\{c_{1},c_{2},\ldots,c_{n}\}$.
Generalized Nash equilibrium is a well-known concept that describes possible outcomes in noncooperative settings. It is a point $\mathbf{x}^{\star}$ where no agent has an incentive to unilaterally change their action, while adhering to constraints that are dependent on other agents’ strategies.
Definition 1 (Generalized Nash equilibrium).
A set of strategies $\mathbf{x}^{\star}=\{x_{1}^{\star},x_{2}^{\star},\ldots,x_{n}^{\star}\}$ is a generalized Nash equilibrium if the following holds for each agent $i\in[n]$,
$c_{i}(x_{i}^{\star},\mathbf{x}_{\neg i}^{\star})\leq c_{i}(x_{i},\mathbf{x}_{\neg i}^{\star}),~{}\forall x_{i}\in\mathbb{X}_{i}(\mathbf{x}_{\neg i}^{\star}).$ (1)
At a generalized Nash equilibrium, each agent’s strategy is optimal given the strategies of all other agents and satisfies the feasibility constraints imposed by the other agents’ actions.
There exist multiple Nash equilibria in many games, meaning that several action profiles $\mathbf{x}^{\star}$ may satisfy the Nash equilibrium condition. This multiplicity arises when different combinations of actions lead to optimal outcomes for each agent. The presence of multiple equilibria introduces a coordination challenge known as the equilibrium selection problem, where agents must agree on one equilibrium from the set of possible options to ensure efficient and safe interaction.
Example 1 (2-agent quadratic game).
Consider two agents, $i$ and $j$, aiming to minimize a shared cost while satisfying a constraint. Each agent’s action $x_{i}\in\mathbb{R}$ is chosen to solve:
$\begin{array}[]{ll}\underset{x_{i}}{\text{minimize}}&c_{i}(x_{i},x_{j})=-(x_{i}+x_{j}-1)^{2},\\ \text{subject to}&x_{i}^{2}+x_{j}^{2}\leq 1.\end{array}$ (2)
A Nash equilibrium $(x_{1}^{\star},x_{2}^{\star})$ satisfies the Karush-Kuhn-Tucker (KKT)conditions. The Lagrangian is:
$\mathcal{L}(x_{1},x_{2},\mu)=-(x_{1}+x_{2}-1)^{2}+\mu(x_{1}^{2}+x_{2}^{2}-1),$ (3)
where $\mu\geq 0$ is the Lagrange multiplier. The first-order conditions are:
$\textstyle{\frac{\partial\mathcal{L}}{\partial x_{1}}=-2(x_{1}+x_{2}-1)+2\mu x_{1}=0,}$ (4a)
$\textstyle{\frac{\partial\mathcal{L}}{\partial x_{2}}=-2(x_{1}+x_{2}-1)+2\mu x_{2}=0,}$ (4b)
$\mu(x_{1}^{2}+x_{2}^{2}-1)=0,\quad\mu\geq 0,\quad x_{1}^{2}+x_{2}^{2}\leq 1.$ (4c)
From the first two equations, we obtain:
${\mu(x_{1}-x_{2})=0,\quad x_{1}+x_{2}=\frac{2}{2-\mu}.}$ (5)
This yields two cases: i) $\mu=0$: The constraint is inactive, The feasible solutions satisfying $x_{1}^{2}+x_{2}^{2}\leq 1$ are all points on the line segment $x_{1}+x_{2}=1$ where $0\leq x_{1},x_{2}\leq 1$. ii) $\mu>0$: The constraint is active, so $x_{1}^{2}+x_{2}^{2}=1$. Since $\mu(x_{1}-x_{2})=0$, we must have $x_{1}=x_{2}$. Solving $12x_{1}^{2}=$ 1 gives $x_{1}=x_{2}=\pm\frac{1}{\sqrt{2}}$.
The critical points are:
${\{(x,1-x)\mid 0\leq x\leq 1\}\cup\left\{\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right),\left(-\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)\right\}.}$ (6)
However, $\{(x,1-x)\mid 0\leq x\leq 1\}$ are local maxima, as an agent can unilaterally decrease their cost by deviating. Thus, $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$ and $\left(-\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)$ are Nash equilibria, demonstrating the existence of multiple equilibria in this game.
In decentralized systems, achieving consensus on which equilibrium to select is nontrivial, especially in noncooperative settings where agents have different preferences over equilibria. Without coordination, there is a risk that agents may settle on different equilibria, resulting in inefficiencies or even safety risks in the system. Furthermore, agents may not be able to communicate directly or share their preferences, further complicating the equilibrium selection problem.
4 Trading auction for consensus
The Trading Auction for Consensus (TACo) algorithm is designed to address the challenge of decentralized conflicting choice problems among non-cooperative agents. The primary goal of TACo is to help agents reach a consensus or agree on a single choice, even when they have conflicting preferences. The algorithm accomplishes this without requiring direct negotiation or communication between agents, and it preserves each agent’s private information during the process.
4.1 Key elements of TACo
TACo consists of several key components: a pay matrix, an offer matrix, a cost matrix, a profit matrix, the private valuations over secondary assets, the amount of trading units, and the decrement factor. These components are central to how agents compute their profit and make decisions during TACo. Let $n\in\mathbb{Z}_{\geq 0}$ denote the number of agents, and $m\in\mathbb{Z}_{\geq 0}$ the number of choices available in the auction. For any matrix $X$, $X_{ij}$ refers to its $(i,j)$ entry.
  1. 1)
    Offer matrix ($\mathcal{O}$): The offer matrix $\mathcal{O}\in\mathbb{R}^{n\times m}$ represents the units that each agent $i$ can receive if they select choice $j$. Specifically, $\mathcal{O}_{ij}$ denotes the number of units that agent $i$ will receive when they choose choice $j$, determined by the offers made by other agents.
  2. 2)
    Pay matrix ($\mathcal{P}$): The pay matrix $\mathcal{P}\in\mathbb{R}^{n\times m}$ captures the cost each agent incurs when selecting a particular choice. $\mathcal{P}_{ij}$ is the number of units that agent $i$ has to pay to others when they choose choice $j$.
  3. 3)
    Cost matrix ($\mathcal{C}$): The cost matrix $\mathcal{C}\in\mathbb{R}^{n\times m}$ reflects the intrinsic cost associated with each agent’s choice. $\mathcal{C}_{ij}$ represents the original cost experienced by agent $i$ when they select choice $j$.
  4. 4)
    Private valuation ($\mathbf{b}$): Each agent $i$ assigns a private valuation, $b_{i}$, to the units of the asset being traded in the TACo algorithm. This valuation, represented as the $i$-th element of the vector $\mathbf{b}\in\mathbb{R}_{>0}^{n}$, reflects how much agent $i$ values each unit of the secondary asset, such as carbon emission credits. The value $b_{i}$ is unique to each agent and remains private, influencing how the agent perceives the offer and pay quantities when calculating their profit.
  5. 5)
    Amount of trading units ($d$): The constant $d\in\mathbb{R}_{>0}$ represents the amount of secondary assets traded in each auction step. This public quantity directly impacts the updates of the offer and pay matrices, influencing each agent’s profit calculation.
  6. 6)
    Trading unit decrement factor ($\gamma$): The decrement factor $\gamma\in(0,1)$ is a parameter used to reduce the number of trading units when a specified condition is met. When triggered, the amount of trading units $d$ is scaled by the decrement factor:
    $d\leftarrow\gamma d.$ (7)
    This reduction adjusts the scale of assets exchanged in the auction, allowing for a controlled decrease in trading volume over time or under certain conditions, which can influence the convergence behavior of the auction process.
  7. 7)
    Profit matrix ($\mathcal{J}$): The profit matrix $\mathcal{J}\in\mathbb{R}^{n\times m}$ represents the profit each agent gains based on their choice. Specifically, $\mathcal{J}_{ij}$ denotes the profit experienced by agent $i$ when selecting choice $j$. The profit is calculated as follows:
    $\mathcal{J}=\text{diag}(\mathbf{b})\cdot(\mathcal{O}-\mathcal{P})-\mathcal{C}.$ (8)
    This equation accounts for each agent’s private valuation b, the offers received $\mathcal{O}$, the payments made $\mathcal{P}$, and the intrinsic costs $\mathcal{C}$, resulting in the total profit for each possible choice.
4.2 Auction mechanism
TACo operates through sequential steps, where agents take turns making decisions. At each step, an agent selects the choice they value most and attempts to persuade others by offering payments. These decisions are based on the matrices $\mathcal{O}$, $\mathcal{P}$, and $\mathcal{C}$. Before explaining the TACo mechanism, we introduce the following definitions.
Definition 2 (Agent-profit matrix tuple $(i,\mathcal{J})$).
The agent-profit matrix tuple $(i,\mathcal{J})$ represents the association of agent index $i\in\mathbb{N}$ with a specific profit matrix $\mathcal{J}$ applied in the current step.
A cycle refers to a sequence of auction processes that returns to a previously encountered profit matrix–playing agent pair. Once the process revisits an existing profit matrix-playing agent pair, the sequence repeats, forming a loop. An interesting property of a cycle is that it includes a state where all agents have reached an agreement, particularly in the trivial case where they continuously select the same choice without termination.
Definition 3 (Cycle, $\sigma$).
A cycle, $\sigma$, is a finite sequence of agent-profit matrix tuples,
$\sigma=\left((i^{k_{1}},\mathcal{J}^{k_{1}}),(i^{k_{2}},\mathcal{J}^{k_{2}}),\ldots,(i^{k_{\ell}},\mathcal{J}^{k_{\ell}})\right),$
in which a tuple $(i^{k_{\ell+1}},\mathcal{J}^{k_{\ell+1}})$ revisits any previously encountered tuple, i.e.,
$(i^{k_{\ell+1}},\mathcal{J}^{k_{\ell+1}})=(i^{k_{1}},\mathcal{J}^{k_{1}}).$
Here, $\ell$ is the length of the cycle.
Definition 4 (Step, $k$).
A step in the TACo process refers to a single update to the profit matrix $\mathcal{J}$, which occurs based on the choice made by one agent during that auction step.
Algorithm 1 Trading Auction for Consensus (TACo)
1:$n$: number of agents, $m$: number of choices, $\epsilon$: termination tolerance
2:$\mathcal{C}\in\mathbb{R}^{n\times m}$, $d_{0}\in\mathbb{R}_{>0}$, $\gamma\in(0,1)$
3:Initialize offer matrix $\mathcal{O}\in\mathbb{N}^{n\times m}$ to $\mathbf{0}_{n\times m}$
4:Initialize pay matrix $\mathcal{P}\in\mathbb{N}^{n\times m}$ to $\mathbf{0}_{n\times m}$
5:selections $\leftarrow$ empty list $\in\mathbb{N}^{n}$
6:isConverged $\leftarrow$ False
7:$d\leftarrow d_{0}$triangleright Initialize trading unit
8:recordedStates $\leftarrow$ empty set $\triangleright$ For cycle detection
9:while not isConverged do
10:     Choose a playing agent $i$ in a sequential manner.
11:     for each choice $j=1$ to $m$ do
12:         Compute profit for agent $i$ and choice $j$:
13:         $\mathcal{J}_{ij}\leftarrow b_{i}(\mathcal{O}_{ij}-\mathcal{P}_{ij})-\mathcal{C}_{ij}$
14:     end for
15:     $j^{\star}\leftarrow\underset{j\in[m]}{\arg\max}(\mathcal{J}_{ij})$triangleright Agent $i$ selects the choice $j^{\star}$ that maximizes its profit
16:     Update the matrices based on the choice $j^{\star}$:
17:      $\mathcal{P}_{ij^{\star}}\leftarrow\mathcal{P}_{ij^{\star}}+nd$
18:      $\mathcal{O}_{ij^{\star}}\leftarrow\mathcal{O}_{ij^{\star}}+d,\forall i\in[n]$
19:     $\text{{selections}}_{i}\leftarrow j^{\star}$triangleright Record agent $i$’s choice
20:     if current $((\mathcal{O}-\mathcal{P}),i)$ exists in recordedStates then $\triangleright$ Detect cycle
21:         Reduce trading unit: $d\leftarrow\gamma\cdot d$
22:         Clear recordedStates
23:         for each agent $i\in[n]$ do
24:              $\max_{j}\mathcal{J}_{ij}-\min_{j}\mathcal{J}_{ij}<\epsilon,\,\forall j\in\text{cycle}$triangleright Check termination condition
25:         end for
26:         if condition is met for all agents then
27:              isConverged $\leftarrow$ True
28:         end if
29:     end if
30:     Add current $((\mathcal{O}-\mathcal{P}),i)$ to recordedStates
31:end while
32:return $\text{mode}(\textit{selections})$triangleright Return the most frequent value in selections
The following rules define the TACo process:
  1.  1)
    Sequential play: Agents take turns sequentially in a predefined order. For example, if there are three agents, the order of play would be agent 1 $\rightarrow$ agent 2 $\rightarrow$ agent 3, and then the sequence repeats.
  2.  2)
    Profit calculation: During agent $i$’s turn, they select the choice $j$ that maximizes their profit, given by:
    $\mathcal{J}_{ij}=b_{i}(\mathcal{O}_{ij}-\mathcal{P}_{ij})-\mathcal{C}_{ij},$ (9)
    where the $i$-th row of each matrix corresponds to the information relevant to agent $i$.
  3.  3)
    Matrix update: When an agent $i$ selects a particular option $j$, the matrices are updated as follows:
    • The algorithm updates the payment matrix $\mathcal{P}$ by incrementing the entry $\mathcal{P}_{ij}$ by $nd$, where $n$ is the total number of agents and $d$ is the trading unit. This reflects that agent $i$ pays $n$ units of the secondary asset for selecting option $j$:
      $\mathcal{P}_{ij}\leftarrow\mathcal{P}_{ij}+nd.$ (10)
    • The algorithm updates the offer matrix $\mathcal{O}$ by incrementing the $j$-th column by $d$. This ensures that all agents receive an additional unit of the secondary asset in their offers for option $j$:
      $\mathcal{O}_{ij}\leftarrow\mathcal{O}_{ij}+d,\quad\forall i\in[n].$ (11)
  4.  4)
    Cycle Detection and Trading Unit Reduction: The trading unit $d$ is reduced by the decrement factor $\gamma$ when a cycle is detected. A cycle is identified by checking whether the same pair of profit matrix $\mathcal{J}$ and active agent at a specific step has occurred in any previous steps. Tracking $\mathcal{J}$ is equivalent to tracking $(\mathcal{O}-\mathcal{P})$, since $b_{i}$ and $C_{ij}$ are constants based on eq. 9. Agents can independently detect cycles by monitoring changes in the publicly observable matrices $\mathcal{O}$ and $\mathcal{P}$. When a cycle is detected, the trading unit is updated as $d_{r+1}\leftarrow\gamma d_{r}$, where $d_{r}$ denotes the trading unit after $r$ cycle detections, with the initial trading unit defined as $d_{0}$. Additionally, the current record of profit matrices is cleared to restart the cycle detection process.
    This reduction of the trading unit is a common practice in real bargaining scenarios, where participants tend to adjust their bids or pricing in smaller increments over time to facilitate reaching a consensus [38]. The relationship governing the trading unit after $r$ cycle detections can be expressed as:
    $d_{r}=d_{0}\cdot\gamma^{r}.$ (12)
  5.  5)
    $\epsilon$-termination: The auction continues until consensus is reached. The process terminates when the difference in profit between all agents’ choice options is within a fixed tolerance $\epsilon$, indicating indifference among the choices in the cycle.
Table 1: TACo running example
Step Agent, $i$ $\mathcal{O}$ $\mathcal{P}$ $\mathcal{J}$ Selections
1 1 $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}-10&-4\\ -7&-9\end{bmatrix}$ $\begin{bmatrix}\underline{2}&\emptyset\end{bmatrix}$
2 2 $\begin{bmatrix}0&1\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&2\\ 0&0\end{bmatrix}$ $\begin{bmatrix}-10&-4.8\\ -7&-7.8\end{bmatrix}$ $\begin{bmatrix}2&\underline{1}\end{bmatrix}$
3 1 $\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$ $\begin{bmatrix}0&2\\ 2&0\end{bmatrix}$ $\begin{bmatrix}\bf{-9.2}&\bf{-4.8}\\ \bf{-8.2}&\bf{-7.8}\end{bmatrix}$ $\begin{bmatrix}\underline{2}&1\end{bmatrix}$
4 2 $\begin{bmatrix}1&2\\ 1&2\end{bmatrix}$ $\begin{bmatrix}0&4\\ 2&0\end{bmatrix}$ $\begin{bmatrix}-9.2&-5.6\\ -8.2&-6.6\end{bmatrix}$ $\begin{bmatrix}2&\underline{2}\end{bmatrix}$
5 1 $\begin{bmatrix}1&3\\ 1&3\end{bmatrix}$ $\begin{bmatrix}0&4\\ 2&2\end{bmatrix}$ $\begin{bmatrix}\bf{-9.2}&\bf{-4.8}\\ \bf{-8.2}&\bf{-7.8}\end{bmatrix}$ $\begin{bmatrix}\underline{2}&2\end{bmatrix}$
Example 2 (TACo running example).
Consider a case with two agents ($n=2$) and two choices ($m=2$). The initial cost matrix is given by $\,\mathcal{C}=\begin{bmatrix}10&4\\ 7&9\end{bmatrix}$, and the agents’ private valuations for the traded asset are $\mathbf{b}=\begin{bmatrix}0.8&1.2\end{bmatrix}$. At each step, agents update their selections based on their calculated profit using the TACo rules. Table 1 shows how the auction progresses over time. In this example, Agent 1 initially selects option 2 because it provides the highest profit. In step 2, Agent 2 selects option 1, which maximizes its own profit. At each step, the agent currently playing updates its selection, and this updated selection is underlined in the table. The bold values in the profit matrix $\mathcal{J}$ highlight that the same profit-value and agent pair reappear at steps 3 and 5, signaling the occurrence of a cycle. This cycle is associated with selection 2, where all agents converge to the same option. The algorithm detects that the $\epsilon$-termination criterion is satisfied because the profit differences for all agents are zero, which is smaller than any positive $\epsilon$. Consequently, the algorithm terminates at step 5.
4.3 Properties of TACo
There are several properties of the TACo procedure (Algorithm 1).
4.3.1 Procedural rationality
TACo incentivizes agents to act in their self-interest by enabling them to trade and negotiate based on their individual valuations of the available options. As demonstrated in Example 2, each agent selects the option that maximizes its profit, naturally driving the auction process toward a consensus.
4.3.2 Privacy preservation
One of TACo’s core strengths is its ability to preserve the privacy of each agent’s valuations. The trading mechanism functions without requiring agents to disclose their private valuations for the available options ($\mathcal{C}$) or trading assets ($\mathbf{b}$).
4.3.3 Independence from direct communication
TACo operates under a minimal communication framework, where agents only need to broadcast their choices. This eliminates the need for direct, 1-on-1 communication, making the algorithm straightforward to implement in practical systems.
4.3.4 Termination guarantee
TACo is guaranteed to terminate with an agreement within a bounded number of rounds. The proof of termination, as well as an explicit upper bound on the number of rounds needed to terminate, is provided in the following section.
5 Termination Proof of TACo
The proof of TACo’s termination (with an agreement) involves three steps. First, we show that TACo always enters a cycle, where the auction steps repeat in a loop as defined in Definition 3. Next, we demonstrate that profit differences between choices within a cycle are bounded, with the bound proportional to the trading unit per step $d$. This restricts each agent’s available choices to options with profit differences no greater than the bound. Finally, we prove that TACo satisfies the epsilon-termination condition and hence converges.
The intuition behind the convergence of TACo lies in that, as soon as a cycle is detected, the trading unit $d$ is reduced by a fixed decrement factor $\gamma$. However, as $d$ decreases, the profit differences within the cycle shrink and must thus eventually fall below the threshold $\epsilon$ for all agents. At this point, the agents reach a consensus by definition, and TACo terminates.
[…]
Figure 2: (Illustration of Lemma 1, $n=4$) This figure states that the row sum $S_{i}^{k}$ of the $i$-th row of the profit matrix $\mathcal{J}$ remains constant every $n$ steps. The fluctuations in the row sum are bounded by the difference $S_{i}^{\text{max}}-S_{i}^{\text{min}}=d(n-1)b_{i}$, where $S_{i}^{\text{max}}$ and $S_{i}^{\text{min}}$ denote the maximum and minimum row sums during the cycle. The diagram highlights the periodic nature of the row sum across steps and the incremental fluctuations $db_{i}$ within each step.
5.1 Existence of cycles in TACo
We first prove that TACo always enters a cycle. A cycle occurs when the auction steps repeat in a loop, as defined in Definition 3. This result is established by analyzing bounded row sums and discrete profit increments. The following lemma shows that the row sum of the profit matrix fluctuates within a bounded range as shown in Figure 2.
Lemma 1 (Row sum of the profit matrix).
The row sum $S_{i}^{k}=\sum_{j=1}^{m}\mathcal{J}_{i,j}^{k}$, where $m$ is the number of choices, remains constant every $n$ steps, with fluctuations equal to:
$S_{i}^{\mathrm{max}}-S_{i}^{\mathrm{min}}=d(n-1)b_{i},$ (13)
where $S_{i}^{\mathrm{max}}$ and $S_{i}^{\mathrm{min}}$ are the maximum and minimum row sums that can occur during the TACo process.
The bounded row sums imply that the actual profit values $\mathcal{J}_{ij}$ are lower bounded, as established in the next lemma.
Lemma 2 (Lower bound of profit values).
The profit value $\mathcal{J}_{ij}$ for any agent $i$ and choice $j$ is lower bounded as
$\mathcal{J}_{ij}\geq\min\left(\min(\mathcal{J}_{i,:}^{0}),\textstyle{\frac{S_{i}^{\mathrm{min}}}{m}}-d(n-1)b_{i}\right),$ (14)
where $m$ is the number of choices and $\min(\mathcal{J}_{i,:}^{0})$ is the minimum element for row $i$ in the initial profit matrix $\mathcal{J}$.
Since the row sums of the profit matrix $\mathcal{J}$ remain constant over every $n$ steps (Lemma 1) and since $\mathcal{J}$ is lower-bounded (Lemma 2), it directly follows that $\mathcal{J}$ is also upper-bounded. We state this consequence in the following corollary.
Corollary 1 (Upper bound of profit values).
The profit values in the matrix $\mathcal{J}$ are upper bounded.
Exploiting the proved upper and lower bounds on $\mathcal{J}$ derived so far, we can now formally prove that TACo always enters a cycle. This result is important because, by definition, a state of consensus where all agents have similar preferences – up to a level of tolerance $\epsilon$ – is also a cycle. Hence, the convergence of TACo will subsequently boil down to showing that the cycle it eventually falls into corresponds to a consensus.
Theorem 1 (The auction always falls into a cycle).
Let $\mathcal{S}$ be the set of all possible agent-profit matrix tuples $(i,\mathcal{J})$, where $i\in[n]$ and $\mathcal{J}$ is a profit matrix satisfying the constraints defined in Lemma 1, Lemma 2 and Corollary 1. The TACo algorithm must eventually revisit a previously encountered tuple $(i,\mathcal{J})\in\mathcal{S}$, resulting in the formation of a cycle.
5.2 Bounded profit differences between choices in a cycle
Having established that TACo always enters a cycle, we now focus on bounding the profit differences among the choices within that cycle. More importantly, we will show that the underlying bound in profit differences is proportional to $d$; a property that will allow us to prove TACo eventually converges to a state of agreement between agents, particularly since $d$ is decremented according to (12) once a cycle is detected. To establish this property, we start by introducing the choice count matrix ($\Delta^{\sigma}$), which tracks how many times each agent selects each choice during a cycle.
Definition 5 (Choice count matrix, $\Delta^{\sigma}$).
The choice count matrix, denoted as $\Delta^{\sigma}\in\mathbb{Z}_{\geq 0}^{n\times m}$, represents the number of times each choice is selected by agents during a cycle. Specifically, $\Delta^{\sigma}_{i,j}$ denotes the number of times agent $i$ selects choice $j$ within the cycle $\sigma$.
It follows that the matrix $\Delta^{\sigma}$ satisfies the following:
$\mathcal{P}^{t+\ell}-\mathcal{P}^{t}=nd\cdot\Delta^{\sigma},$ (15)
$\mathcal{O}^{t+\ell}-\mathcal{O}^{t}=d\cdot H_{n}\Delta^{\sigma},$ (16)
where $\mathcal{P}^{t+\ell}-\mathcal{P}^{t}$ and $\mathcal{O}^{t+\ell}-\mathcal{O}^{t}$ represent the net changes in the pay matrix and offer matrix after each cycle of length $\ell$, respectively, and $H_{n}=\textbf{1}_{n}\textbf{1}_{n}^{\top}$ is an $n\times n$ matrix with all elements equal to 1, capturing the aggregation of offers received across all agents. Using the definition of the choice count matrix and its relationship shown in eq. 15 and eq. 16, we establish the structure of $\Delta^{\sigma}$ which shows that, within a cycle, all agents select each choice the same number of times.
Lemma 3 (Uniform choice count matrix in a cycle).
The choice count matrix $\Delta^{\sigma}\in\mathbb{Z}_{\geq 0}^{n\times m}$ for a cycle has the form
$\Delta^{\sigma}=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{m}\\ \vdots&\vdots&\ddots&\vdots\\ c_{1}&c_{2}&\ldots&c_{m}\end{bmatrix},$ (17)
where each $c_{j}\in\mathbb{Z}_{\geq 0}$ represents the number of times that choice $j$ is selected by each agent within the cycle.
This uniform structure naturally leads to the concept of active choices: the choices selected by agents during a cycle.
Definition 6 (Active choices in a cycle, $A^{\sigma}$).
The active choices in a cycle, denoted $A^{\sigma}$, are the row indices of the choice count matrix $\Delta^{\sigma}$ with non-zero values. In other words, $A^{\sigma}$ is a set of indices representing the choices selected at least once by some agent within a cycle.
We now analyze the profit differences within a cycle by focusing on active choices only. Leveraging Lemma 1, Lemma 2, and Lemma 3, we prove a key result that bounds these profit differences proportionally to the amount of trading units $d$.
Theorem 2 (Bound on profit differences within a cycle).
The profit difference among the active choices $A^{\sigma}$ in a cycle $\sigma$ for agent $i$ is bounded as
$U_{\sigma}^{i}-L_{\sigma}^{i}\leq(m+1)d(n-1)b_{\max},\quad\forall i\in[n],$ (18)
where $m$ is the total number of choices, $d$ is the trading unit for the current cycle, $b_{\max}=\max_{i\in[n]}(b_{i})$ is the highest private valuation among the agents, and $U_{\sigma}^{i}$ and $L_{\sigma}^{i}$ are the maximum and minimum profit values for the active choices of agent $i$ within cycle $\sigma$, respectively.
5.3 Finite termination property of TACo
With the profit differences bounded proportionally to $d$ (Theorem 2), we now demonstrate that TACo satisfies the epsilon-termination condition, leading to an approximate agreement in a finite number of steps. Recall that each time a cycle is detected – which provably happens according to Theorem 1 – the algorithm reduces the trading unit $d$ by a fixed decrement factor $\gamma$. This reduction ensures that the profit differences between choices eventually fall below the threshold $\epsilon$, leading to termination.
Theorem 3 (Termination of TACo).
The TACo algorithm terminates in a finite number of steps.
We also provide a practical bound on the worst-case number of steps required till termination. This bound accounts for the number of cycles needed to meet the epsilon-termination condition as well as the steps within each cycle, hence providing practical insights into the efficiency of TACo.
Theorem 4 (Termination bound for TACo).
The TACo algorithm terminates at most within the following number of steps:
${\left\lceil\log_{\gamma}\left(\frac{\epsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil\cdot n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}},$ (19)
with decrement factor $\gamma$ and initial trading unit $d_{0}$.
Theorem 4 implies that if $\epsilon$ is increased to $\frac{\epsilon}{\gamma}$, then the worst-case termination bound decreases by $n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}$ steps.
6 Numerical Experiments
6.1 Waypoint merging game
We model a common noncooperative scenario in air traffic management, where $n$ aircraft converging on a waypoint must adjust their speeds to avoid collisions, as shown in Figure 1. Each aircraft $i$ in this waypoint merging game has an estimated time of arrival (ETA) $e_{i}$ and seeks to minimize deviations from it while ensuring safe separation. We introduce carbon emission credits as a tradable secondary asset.
Each aircraft (agent) $i$ selects an action $x_{i}$, representing its adjustment to the ETA. The adjusted arrival time for agent $i$ is then $u_{i}=e_{i}+x_{i}$. The cost structure for each agent is:
$\begin{array}[]{ll}\underset{x_{i}}{\mbox{minimize}}&c_{i}(x_{i})=k_{i}x_{i}^{2}\\ \mbox{subject to}&|u_{i}-u_{j}|\geq D,\enskip\forall j\in[n]\setminus i,\end{array}$ (20)
where $k_{i}>0$ represents the urgency factor, indicating sensitivity to arrival time adjustments, and $D\geq 0$ is the required separation distance.
This game Equation 20 admits multiple generalized Nash equilibria. The absolute value in the constraint introduces two conditions depending on the arrival sequence. Specifically, the sign of the terms inside the absolute value changes based on the order of $u_{i}$ and $u_{j}$, leading to two inverted inequalities. This leads to $n!$ distinct equilibria. Thus, coordination is essential for safe and efficient merging.
Another key challenge is the lack of direct 1-on-1 communication between aircraft. Instead, aircraft use the Automatic Dependent Surveillance-Broadcast (ADS-B)system, which broadcasts publicly available information such as positions, urgency factors ($k_{i}$), and ETAs ($e_{i}$). However, ADS-B does not support direct negotiation, making consensus on an equilibrium difficult despite agents being able to compute all equilibria.
In our scenario, each aircraft privately determines its valuation for emission credits and employs the TACo algorithm to iteratively negotiate arrival adjustments. By trading emission credits and broadcasting preferences via ADS-B, aircraft reach a consensus while acting in their own self-interest, despite the communication constraints imposed by ADS-B.
6.2 Baseline algorithms
We evaluated the effectiveness of the TACo algorithm by comparing it with widely used negotiation approaches. While these baseline methods lack key features of TACo, they serve as intuitive benchmarks for multi-agent coordination. The baseline algorithms considered in this experiment are:
  • Voting: Each agent votes for their preferred options, and the final decision is determined by majority rule [39]. In the event of a tie, the winner is chosen randomly.
  • Random dictator: A random agent is selected, and all other agents must conform to the preferences of this chosen agent.
  • Utilitarian: A central coordinator selects the option that minimizes the total costs across all agents, requiring compliance from every agent.
  • Egalitarian: A central coordinator selects the option that maximizes fairness among agents, and all agents are required to comply with the decision.
The voting mechanism offers a simple, decentralized approach but can lead outcomes which are not unilaterally rational, as agents must accept majority decisions even when misaligned with their preferences. Moreover, the utilitarian and egalitarian methods optimize system-level objectives but impose centralized decisions, making them unsuitable for noncooperative and decentralized settings. Lastly, the random dictator mechanism also relies on centralized selection but ensures fairness over time in repeated interactions by rotating decision-making power among agents.
6.3 Evaluation method
We evaluated the performance of TACo across three key metrics: optimality gap, fairness, and convergence steps.
  • Optimality gap: The optimality gap quantifies the deviation of the total cost incurred by the algorithm from the optimal total cost achieved by a utilitarian solution. It is defined as:
    ${\mathrm{OG}(\mathbf{c},\mathbf{c}^{\star})=\frac{\sum_{i\in[n]}c_{i}}{\sum_{i\in[n]}c_{i}^{\star}}-1,}$ (21)
    where $c_{i}^{\star}\in\mathbf{c}^{\star}$ is the cost of agent $i$ from the utilitarian solution, and $c_{i}$ is the cost of agent $i$ under the evaluated algorithm.
  • Gini index: The Gini index [40] measures fairness among agents. A smaller Gini index indicates more equitable outcomes. It is computed as:
    ${\mathrm{GI}(\textbf{c})=\frac{1}{2n\sum_{i\in[n]}c_{i}}\sum_{i,j\in[n]}|c_{i}-c_{j}|.}$ (22)
  • Convergence steps: The convergence steps metric measures the algorithm’s efficiency by counting the number of steps required for termination.
To assess the effectiveness of TACo, we conducted a series of Monte Carlo simulations with 1000 trials to ensure statistical significance. Each trial incorporated randomized initial conditions, including diverse weight parameters ($k_{i}$) and varying initial estimated times of arrival. Agents’ private valuations over secondary assets ($b_{i}$) were randomly assigned within a predefined range to reflect diverse private information. The decrement factor $\gamma$ was set to 0.9, with the number of aircraft $n$ set to 4 and the number of choices $m$ set to 24.
The simulations were implemented in the Julia programming language. We used the ParametricMCPs.jl package [41] and the PATH solver for mixed complementarity problems [42] to compute the Nash equilibria of the game Equation 20 prior to running TACo.
6.4 Results
(a) Optimality gap
(b) Gini index
Figure 3: Experimental results comparing the performance of the TACo algorithm with baseline methods (Voting, Utilitarian, Egalitarian, and Random dictator) across three evaluation metrics: optimality gap, Gini index, and termination steps. TACo demonstrates the smallest median values for the optimality gap and Gini index, highlighting its effectiveness in achieving near-optimal and fair outcomes.
6.4.1 Optimality and fairness
As shown in Figure 3(a), the TACo algorithm exhibited a consistently lower optimality gap compared to the baseline algorithms, except for Utilitarian, which achieves a zero optimality gap by definition. TACo showed minimal deviation from the optimal solution, with a median and third quartile value of 0% and a maximum value of 20.3%. This confirms that the TACo effectively minimizes deviations from the socially optimal solution.
Figure 3(b) shows that TACo achieves a low Gini index, with a median of 0.181, demonstrating that the TACo improves fairness by redistributing costs among agents. Notably, trading-based coordination enhances fairness beyond merely selecting the most balanced equilibrium. However, in cases where the cost differences between choices were initially small, the trading process amplified disparities, resulting in occasional outliers. This suggests that the choice of trading unit $d$ should account for the relative magnitude of costs in the system.
6.4.2 Termination proof validation
In our experiments, TACo reached consensus in a median of 53 steps, with a maximum of 380 steps. Given an ADS-B information exchange rate of 1 Hz, this corresponds to a median negotiation time of 53 seconds and a maximum of 6 minutes and 20 seconds. While TACo is iterative by design, these results suggest that its convergence remains within operationally feasible limits.
Figure 4: Observed profit differences between choices in cycles, compared with theoretical upper bounds. The solid blue line represents the maximum upper bound derived from Equation 18 with $m=24$, the maximum number of active choices in this scenario. Since most cycles involved only two choices, a dotted line provides an additional upper bound for $m=2$ for comparison. The results confirm that profit differences consistently remained below the theoretical bounds.
As shown in Figure 4, the maximum observed profit difference between choices remained below the theoretical upper bound from Equation 18, validating that profit differences decrease as predicted. This empirical confirmation supports the theoretical termination guarantee of TACo.
6.5 TACo variations for faster convergence
To accelerate the convergence speed of the TACo algorithm, we explored two alternative strategies: (i) adjusting the decrement factor and (ii) interrupting the algorithm to enforce consensus earlier. These experiments highlight how convergence speed can be improved, albeit with potential trade-offs in rationality, optimality, and fairness.
(a) Termination steps
(b) Optimality gap and Gini index
Figure 5: Effect of the decrement factor on TACo performance. (a) The termination steps increase as the decrement factor $\gamma$ approaches 1, with a maximum exceeding 1500 steps. For $\gamma\leq 0.9$, there is no notable reduction in convergence steps. (b) The Gini index and the optimality gap remain largely unaffected by changes in $\gamma$, demonstrating that TACo maintains fairness and optimality regardless of the decrement factor.
(a) Termination steps
(b) Optimality gap and Gini index
Figure 6: Effect of interruptions on TACo performance. (a) Interrupting the algorithm earlier reduces termination steps effectively. (b) However, earlier interruptions increase the optimality gap and Gini index. This demonstrates the trade-off between faster convergence and reduced fairness and optimality.
6.5.1 Effect of decrement factor on convergence
We investigated how varying the decrement factor $\gamma$ impacts both convergence steps and performance. In this experiment, we tested $\gamma$ values ranging from 0.3 to 0.99. As shown in Figure 5(a), the highest number of steps occurred in the $\gamma=0.99$ case, with a median of 57 and a maximum of 1680 steps. However, for $\gamma$ values below 0.9, there was no significant improvement in the convergence rate, even for the $\gamma=0.3$ case.
A similar trend was observed in the Gini index and optimality gap, as shown in Figure 5(b). Just as there was no significant improvement in the convergence rate for smaller $\gamma$ values, there was also no notable impact on optimality or fairness. Thus, as long as $\gamma$ does not exceed 0.9, there is no significant impact on convergence steps, optimality, or fairness.
6.5.2 Effect of algorithm interruption on convergence
In this experiment, we interrupted the algorithm before reaching natural consensus to simulate scenarios where agents must reach a decision within a limited time. Agents were forced to select the most commonly chosen option up to the interrupted step, sacrificing individual rationality for the sake of faster decision-making.
Figure 6 shows the results of these interruptions, comparing termination steps and performance across varying levels of interruption. The results demonstrated that interruptions significantly reduced convergence steps, as shown in Figure 6(a). However, as expected, forced consensus increased both the optimality gap and Gini index, as seen in Figure 6(b).
In the most extreme case, where the algorithm was interrupted at the fifth step—allowing each agent only a single chance to express their preference—the median optimality gap increased from 0% to 182.7%, and the median Gini index rose from 0.181 to 0.449. This experiment demonstrates that TACo can be adapted to systems requiring faster convergence by allowing interruptions, albeit at the cost of reduced optimality, fairness, and rationality. Such adaptations may be appropriate in high-stakes environments where timely decision-making is more critical than achieving perfect rationality.
7 Conclusion
We developed the trading auction for consensus (TACo) algorithm, a decentralized approach for coordinating decisions among noncooperative agents facing conflicting choices. TACo enables agents to reach consensus by trading secondary assets, ensuring procedural rationality without requiring centralized coordination, direct communication, or the disclosure of private information. This flexibility makes TACo an effective tool for resolving coordination challenges in multi-agent systems.
Future work includes improving scalability, evaluating performance in more complex environments, and tightening convergence bounds to ensure TACo terminates quickly in safety-critical settings. This will clarify the types of problems it can efficiently solve and inform strategies for faster convergence.
Appendix: Proofs of Main Results
Proof of Lemma 1.  Over every $n$ steps: 1. During its turn, agent $i$ updates $\mathcal{P}$ according to eq. 10, resulting in a payment of $d(n-1)b_{i}$ to other agents. 2. During the $n-1$ turns of other agents, agent $i$ receives $db_{i}$ from each, by updating the offer matrix $\mathcal{O}$ according to eq. 11.
Thus, the total amount paid and received for agent $i$ balances out over $n$ steps, keeping the row sum $S_{i}^{k}$ constant every $n$ steps. The maximum fluctuation in $S_{i}^{k}$ occurs when agent $i$ makes its payment (reducing the row sum by $d(n-1)b_{i}$), and the row sum is restored during subsequent steps when agent $i$ receives offers.
Proof of Lemma 2.  By Lemma 1, the row sum $S_{i}^{k}$ of the $i$-th row of the profit value matrix is constant every $n$ steps, and the average profit per choice at an arbitrary step $k$ is $\frac{S_{i}^{k}}{m}$. In addition, only choices with a profit above this average can be selected by agent $i$, because the choices agent $i$ selects must maximize profit. Moreover, once a choice is selected, its profit decreases by $(n-1)db_{i}$. Aggregating all this information, the profit value $\mathcal{J}_{ij}$ of choice $j$ for agent $i$ cannot voluntarily drop below:
${\frac{S_{i}^{\mathrm{min}}}{m}}-d(n-1)b_{i}.$ (23)
On the other hand, there might exist choices for agent $i$ with a profit value below (23) due to the way they were initialized. Combining this fact with the (23), it follows that
$\mathcal{J}_{ij}\geq\min\left(\min(\mathcal{J}_{i,:}^{0}),\ \textstyle{\frac{S_{i}^{\mathrm{min}}}{m}}-d(n-1)b_{i}\right).$ (24)
Proof of Theorem 1.  By Lemma 2 and Corollary 1, each row of the profit matrix $\mathcal{J}$ is bounded above and below, with bounds denoted by $U$ and $L$, respectively. According to Algorithm 1, the profits can vary only in discrete steps of size $db_{i}$, where $d>0$ is the trading unit and $b_{i}>0$ is the agent-specific weight. Thus, each element $\mathcal{J}_{ij}$ of $\mathcal{J}$ can assume only a finite number of values, given by:
${\left\lfloor\frac{U-L}{db_{i}}\right\rfloor}.$ (25)
Since the total number of configurations of $\mathcal{J}$ is finite, this means that the set $\mathcal{S}$ of all possible tuples $(i,\mathcal{J})$ is finite as well. As the TACo algorithm iterates, it progresses through elements of this finite set $\mathcal{S}$. By the pigeonhole principle, the TACo algorithm must eventually revisit a previously encountered tuple $(i,\mathcal{J})$ and hence fall into a cycle.
Proof of Lemma 3.  For a cycle to form, the profit matrix $\mathcal{J}$ must return to a previously encountered state. By (9), this requires that the total amount an agent pays for a specific choice exactly balances the total offers received from other agents for that choice. Mathematically, this condition can be expressed as:
$\mathcal{P}^{t+\ell}-\mathcal{P}^{t}=\mathcal{O}^{t+\ell}-\mathcal{O}^{t}.$ (26)
From (15)-(16), (26) is equivalent to
$nd\cdot\Delta^{\sigma}=d\cdot H_{n}\Delta^{\sigma},$ (27)
from which we obtain
$(n\textbf{I}_{n}-H_{n})\Delta^{\sigma}=0,$ (28)
where $\textbf{I}_{n}$ is the $n\times n$ identity matrix.
Note that in (28), the matrix $(n\textbf{I}_{n}-H_{n})$ is the Laplacian matrix of a fully connected graph with $n$ nodes. As shown in [43, p.8], for such a Laplacian one must have $\det(n\textbf{I}_{n}-H_{n})=0$, which indicates the existence of nontrivial solutions to (28). Transforming $(n\textbf{I}_{n}-H_{n})$ into row echelon form using Gaussian elimination yields:
$n\cdot\underbrace{\begin{bmatrix}\textbf{I}_{n-1}&-\textbf{1}_{n-1}\\ \textbf{0}_{n-1}^{\top}&0\end{bmatrix}}_{A}\Delta^{\sigma}=0.$ (29)
From this structure, the first row of $A$ implies that the first and last elements of each column of $\Delta^{\sigma}$ are identical. The second row of $A$ implies that the second and last elements of each column of $\Delta^{\sigma}$ are also identical. By induction, all elements within each column of $\Delta^{\sigma}$ are identical. Therefore, $\Delta^{\sigma}$ must have the form:
$\Delta^{\sigma}=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{m}\\ \vdots&\vdots&\ddots&\vdots\\ c_{1}&c_{2}&\ldots&c_{m}\end{bmatrix},$ (30)
where $c_{j}\in\mathbb{Z}_{\geq 0}$ for all $j\in[m]$. This structure implies that, for a cycle to form, each agent must select each choice the same number of times as every other agent within the cycle. This completes the proof.
Proof of Theorem 2.  From Theorem 1, TACo must eventually fall into a cycle $\sigma$. Let the number of active choices within that cycle be $p=|A^{\sigma}|$, and define $C_{i}$ as the sum of the profits for row $i$ corresponding to the inactive choices, i.e., the choices not in the cycle $\sigma$. Within $\sigma$, the sum of profits for the active choices in row $i$ is:
$S_{i}^{k}-C_{i},$ (31)
where $C_{i}$ remains constant during the cycle. Following the same logic as in Lemma 2, the lower bound for the profits of the active choices in $\sigma$ becomes:
$\mathcal{J}_{ij}\geq\min\left(\min(\mathcal{J}_{i,:}^{\text{0}}),\textstyle{\frac{S_{i}^{\mathrm{min}}-C_{i}}{p}}-d(n-1)b_{i}\right),\quad\forall j\in A^{\sigma}.$ (32)
From Lemma 3, each active choice $j\in A^{\sigma}$ is selected at least once by each agent. For such a choice to be selected, it must, at the time of selection, have a profit value greater than the average profit $\frac{S_{i}^{\mathrm{min}}-C_{i}}{p}$ among the active choices in $\sigma$. Accordingly, no choice that has a profit value below $\frac{S_{i}^{\mathrm{min}}-C_{i}}{p}$ may be selected by agent $i$. Therefore, strictly within the active choices in $\sigma$, we obtain the following tighter lower bound on $\mathcal{J}$:
$\mathcal{J}_{ij}\geq\textstyle{\frac{S_{i}^{\mathrm{min}}-C_{i}}{p}}-d(n-1)b_{i}=L_{\sigma}^{i},\quad\forall j\in A^{\sigma}.$ (33)
Since the row sum of the profit matrix is constant over every $n$ steps, this lower bound ensures the existence of an upper bound for the profits of the active choices. The upper bound, $U_{\sigma}^{i}$, occurs when one active choice achieves a profit value above the average, while the remaining $p-1$ active choices are at their lower bound. This situation leads to the following relation:
$L_{\sigma}^{i}(p-1)+U_{\sigma}^{i}=S_{i}^{\mathrm{max}}-C_{i}.$ (34)
Rearranging for $U_{\sigma}^{i}-L_{\sigma}^{i}$ yields
$U_{\sigma}^{i}-L_{\sigma}^{i}=S_{i}^{\mathrm{max}}-C_{i}-L_{\sigma}^{i}p.$ (35)
Substituting $L_{\sigma}^{i}$ from eq. 33 in (35),
$U_{\sigma}^{i}-L_{\sigma}^{i}$ $=\textstyle{S_{i}^{\mathrm{max}}-C_{i}-p\left(\frac{S_{i}^{\mathrm{min}}-C_{i}}{p}-d(n-1)b_{i}\right)}$ (36)
$=S_{i}^{\mathrm{max}}-C_{i}-S_{i}^{\mathrm{min}}+C_{i}+pd(n-1)b_{i}$ (37)
$=S_{i}^{\mathrm{max}}-S_{i}^{\mathrm{min}}+pd(n-1)b_{i}.$ (38)
From Lemma 1, the row sum difference $S_{i}^{\mathrm{max}}-S_{i}^{\mathrm{min}}$ is equal to $d(n-1)b_{i}$. Substituting this result:
$U_{\sigma}^{i}-L_{\sigma}^{i}=d(n-1)b_{i}+pd(n-1)b_{i}=(p+1)d(n-1)b_{i}.$ (39)
Since $p\leq m$, where $m$ is the total number of choices, the profit difference is bounded as:
$U_{\sigma}^{i}-L_{\sigma}^{i}\leq(m+1)d(n-1)b_{i}\leq(m+1)d(n-1)b_{max},$ (40)
for all $i\in[n]$. This concludes the proof.
Proof of Theorem 3.  By Theorem 1, the TACo process will always fall into a cycle. If the process enters a cycle involving a single choice, all agents consistently select the same choice. In this state, there is no profit difference, and the $\epsilon$-termination criterion is automatically satisfied.
For cycles involving multiple choices, from Theorem 2, the profit difference between any two choices for agent $i$ within a cycle is bounded as
$\max_{(i,\mathcal{J}^{k}),(i,\mathcal{J}^{k^{\prime}})\in\sigma}|\mathcal{J}^{k}_{i,j}-\mathcal{J}^{k^{\prime}}_{i,j^{\prime}}|\leq U_{\sigma}^{i}-L_{\sigma}^{i}\leq(m+1)d(n-1)b_{\max},$ (41)
for all $j,j^{\prime}\in[m]$ and for all $i\in[n]$, where $k$ and $k^{\prime}$ are steps within the cycle. According to the trading unit reduction rule (item 4 in Section 4.2), when a cycle is detected (i.e., when a previously encountered profit matrix $\mathcal{J}$ reappears), the trading unit $d$ is reduced by a decrement factor $\gamma$, where $10<\gamma<$ 1. Each reduction in $d$ proportionally decreases the profit differences across choices, ensuring convergence toward zero.
Specifically, as denoted in Section 4.2, if $d_{r}$ is the trading unit after $r$ cycle detections, then
$d_{r}=d_{0}\cdot\gamma^{r},$ (42)
where $d_{0}$ is the initial trading unit. After $r$ reductions, from eq. 41, the maximum profit difference for any agent is bounded by the value
$(m+1)d_{r}(n-1)b_{\max}=(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}.$ (43)
Since $\gamma<1$, and $m$, $n$, and $b_{\max}$ are constants, the term $(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}$ approaches zero as $r\to\infty$. Hence, for any $\epsilon>0$, there exists a finite $R$ such that for all $r\geq R$:
$\max_{(i,\mathcal{J}^{k}),(i,\mathcal{J}^{k^{\prime}})\in\sigma}|\mathcal{J}^{k}_{i,j}-\mathcal{J}^{k^{\prime}}_{i,j^{\prime}}|\leq\epsilon,\quad\forall i\in[n],\forall j,k\in A^{\sigma}.$ (44)
At this point, the $\epsilon$-termination criterion is satisfied, and the TACo process terminates.
Proof of Theorem 4.  The profit difference among the active choices for agent $i$ is given in Equation 39. Substituting this into Equation 25, we infer that the number of possible values that $\mathcal{J}_{i,j}$ can take in a cycle is at most $(p+1)(n-1)$, where $p$ is the number of active choices.
Since $p\leq m$ and the number of elements in $\mathcal{J}$ is $nm$, the maximum number of distinct $\mathcal{J}$ configurations is $\big{(}(m+1)(n-1)\big{)}^{nm}$. Consequently, the maximum number of unique agent-profit matrix tuples $(i,\mathcal{J})$ in a cycle is $n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}$. This provides an upper bound on the number of steps required to form a cycle, as established in Theorem 1.
The upper bound on the number of cycles $r$ required to satisfy the termination condition is derived from Equation 43 and Equation 44:
$(m+1)d_{0}(n-1)b_{\max}\cdot\gamma^{r}\leq\epsilon,$ (45a)
$\textstyle{\implies r=\left\lceil\log_{\gamma}\left(\frac{\epsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil,}$ (45b)
with decrement factor $\gamma$ and initial trading unit $d_{0}$.
Combining the number of cycles needed and the maximum steps per cycle, we obtain the upper bound on the number of steps needed for TACo to terminate:
${\left\lceil\log_{\gamma}\left(\frac{\epsilon}{(m+1)d_{0}(n-1)b_{\max}}\right)\right\rceil\cdot n\cdot\big{(}(m+1)(n-1)\big{)}^{nm}.}$ (46)
{IEEEbiography}
[[Uncaptioned image]]Jaehan Im (Graduate Student Member, IEEE) was born in Seoul, Republic of Korea. He received the B.S. and M.S. degrees in Aerospace Engineering from the Korea Advanced Institute of Science and Technology (KAIST). He is currently pursuing the Ph.D. degree in the Department of Aerospace Engineering and Engineering Mechanics at the University of Texas at Austin.
{IEEEbiography}
[[Uncaptioned image]]Filippos Fotiadis (Member, IEEE) was born in Thessaloniki, Greece. He received the PhD degree in Aerospace Engineering in 2024, and the MS degrees in Aerospace Engineering and Mathematics in 2022 and 2023, all from Georgia Tech. Prior to his graduate studies, he received a diploma in Electrical & Computer Engineering from the Aristotle University of Thessaloniki. He is currently a postdoctoral researcher at the Oden Institute for Computational Engineering & Sciences at the University of Texas at Austin. His research interests are in the intersection of systems & control theory, game theory, and learning, with applications to the security and resilience of cyber-physical systems.
{IEEEbiography}
[[Uncaptioned image]]Daniel Delahaye (Member, IEEE) is the head of the Optimization and Machine Learning Team of the ENAC research laboratory and he is also in charge of the research program “AI4DECARBO” in the new AI institute ANITI in Toulouse and member of the SESAR scientific committee. He obtained his engineering degree from the ENAC school and did a master of science in signal processing from the National Polytechnic Institute of Toulouse in 1991. He received his PH.D in automatic control from the Aeronautic and Space National School in 1995 under the co-supervision of Marc Schoenauer (CMAPX). He did a post-doc at the Department of Aeronautics and Astronautics at MIT in 1996 under the supervision of Pr Amedeo Odoni. He started his career working at the French Civil Aviation Study Center (CENA) and moved to ENAC in 2008. He got his tenure in applied mathematics in 2012. He conducts research on mathematical optimization and artificial intelligence for airspace design and aircraft trajectory optimization.
{IEEEbiography}
[[Uncaptioned image]]Ufuk Topcu(Fellow, IEEE) is currently a Professor with the Department of Aerospace Engineering and Engineering Mechanics, The University of Texas at Austin, Austin, TX, USA, where he holds the Temple Foundation Endowed Professorship No. 1 Professorship. He is a core Faculty Member with the Oden Institute for Computational Engineering and Sciences and Texas Robotics and the director of the Autonomous Systems Group. His research interests include the theoretical and algorithmic aspects of the design and verification of autonomous systems, typically in the intersection of formal methods, reinforcement learning, and control theory.
{IEEEbiography}
[[Uncaptioned image]]David Fridovich-Keil (Member, IEEE) received the B.S.E. degree in electrical engineering from Princeton University, and the Ph.D. Degree from the University of California, Berkeley. He is an Assistant Professor in the Department of Aerospace Engineering and Engineering Mechanics at the University of Texas at Austin. Fridovich-Keil is the recipient of an NSF Graduate Research Fellowship and an NSF CAREER Award.